Garth has $50,000 to invest for retirement, how much money will he have after 20 years if he buys a CD that earns an 8% APR compounded monthly?
Answer #1

We use the formula:  
A=P(1+r/12)^12n
where
A=future value
P=present value  
r=rate of interest
n=time period.

A=50,000*(1+0.08/12)^(12*20)

=50,000*4.92680277

=$246340.14(Approx)